OpenStudy (mimi_x3):

I'm trying to change this sturm liouville equation into a boundary value problem. To no avail I cannot get it work. Can anyone help me out <3

OpenStudy (mimi_x3):

with \(x = e^t\)

OpenStudy (irishboy123):

.

OpenStudy (mimi_x3):

what does \(.\) mean

OpenStudy (empty):

What did you try?

OpenStudy (mimi_x3):

a bunch of scribbles getting no where I AM LOST

OpenStudy (mimi_x3):

this is fked i swear i should ahve done operating systems

OpenStudy (mimi_x3):

with x = e^t i should get to this https://gyazo.com/2405ad6a1ffabe622aa8c8d0939c65c0

OpenStudy (mimi_x3):

but i am not getting there i am going in circles

OpenStudy (mimi_x3):

i am getting like a humongous bunch of eqns

OpenStudy (michele_laino):

please try this substitution: \(\huge y=kx^s\)

OpenStudy (mimi_x3):

i need this x = e^t

OpenStudy (empty):

You shoulda done linear algebra bb

OpenStudy (empty):

sry u don't know the chain rule I can't help you.

OpenStudy (mimi_x3):

so help me with chain rule

OpenStudy (jadedry):

I'm afraid this is beyond my current level, but if you want to learn about chain rules, you can go here: https://www.khanacademy.org/math/differential-calculus/taking-derivatives/chain_rule/v/chain-rule-introduction Khan Academy always provides free, detailed tutorials on these things. c:

OpenStudy (irishboy123):

getting to that thing you posted is not that hard. forget about sturm-louiville for a bit, that's the bit that's throwing me. i don't know what that reference means in the problem. this is a euler-caucy 2nd order linear DE, and you using the [much better, IMHO] change of independent variable substitution, \(x=e^t\), to solve ot this means \(xy' = Y'\), and \(x^2y'' = Y'' - Y'\) which you sub straight into your DE now in terms of independent variable t.[i presume your uppercase Y signifies that we are now working with our new independent varable t so i gave followed that convention.] so your DE in x becomes in t : \(Y'' - Y' + Y'+ \lambda Y = 0\) or \(Y'' + \lambda Y = 0\) which is what you want. to change the boundary values, again using \(x = e^t\) we have \(y(1)\) = 0, so \( Y(0) = 0\) as \(1 = e^0\), and \(y'(e) = Y'(1) = 0\). as for eigen stuff, i assume you are now going to create a system?

OpenStudy (irishboy123):

and if you can wait, someone like @SithsAndGiggles or @freckles is well worth waiting for.....

OpenStudy (anonymous):

Borrowing from @IrishBoy123's comment, a change of variables \(x=e^t\), \(y(x)=Y(t)\) yields the linear ODE, \[Y''+\lambda Y=0~~\implies~~Y=C_1\cos(\sqrt\lambda t)+C_2\sin(\sqrt\lambda t)\]which is equivalent to \[y=C_1\cos(\sqrt\lambda \ln x)+C_2\sin(\sqrt\lambda \ln x)\]With the given boundary conditions, you get the following system: \[\begin{cases}C_1\cos0+C_2\sin0=0\\[1ex] -C_1\dfrac{\sqrt\lambda}{e}\sin(\sqrt \lambda)+C_2\dfrac{\sqrt\lambda}{e}\cos(\sqrt\lambda)=0\end{cases}\implies\begin{cases}C_1=0\\[1ex]C_2\cos(\sqrt\lambda)=0\end{cases}\]You want a nontrivial solution, so you require that \(C_2\neq0\). It should be easy to find the eigenvalues now.

OpenStudy (anonymous):

Finding the eigenfunctions is easy enough once you determine \(\lambda\), so you would have \(y=\sin(\sqrt\lambda\ln x)\).